Minutes of the Tracker Meeting

October 14, 2004


disclaimer: all time estimates referring to work being done “next week” are probably delayed by the new stand-down and safety training.


Action Items:

  1. Jeff: order ESD caps
  2. Arthur: check on getting prepreg test data from COI
  3. Jeff: procure desiccant bags and ESD wrap


1) Cable-3 problem update:

      - The problem will be investigated further after tower disassembly.  The situation at present is that we know it to be a problem with the connector on the top MCM, not a cable problem.  The MCM was a preproduction one, which has had a lot of use (but always with connector savers attached).  There may have been some damage in the cable installation, perhaps due to the fact that no bending tool was available for that cactus arm, making it difficult to mate and putting a lot of stress onto the connectors.  We have to have the bending tool for Tower A!  Also, there needs to be some stress relief in that area, provided by the sidewall (i.e. foam or O-ring between this cactus arm and the sidewall).

2) Tracker test temperatures & document update & release

      - Nicola Mazziotta will update the tray thermal test plan and procedure to go up to 55C (used to be 50C).  Done.

      - Francesco Belli will update the tray panel T/V test plan to 60C (officially used to be 85C).  Done.

      - Robert will ask Natalie to re-release both.  Done.

3) Bias circuit coupon testing closure plan

      - Dave sent out a new drawing for approval, and Nick has signed off on it.  The notes have been changed such that no via contact is required, and therefore no coupon testing is needed.  Once released, this will recover all of the bias circuits presently on hold.

4) Tall MCMs for Tower-A

      - Robert will hand carry 5 of them to Pisa on Sunday.  Done.

5) Flex-circuit cables

      - Almost all of the original Parlex production has failed one coupon test or another.  This casts just as much suspicion on those that did pass, since the pass could have been fortuitous and does not necessarily mean that the vias in the actual cables are good.

      - All vias tested by GSFC from the new, improved Parlex process were good. 

      - Our hope is to get a new Tower-A set by the end of October, so that we can test them and ship them to Italy just in time for Tower-A assembly. 

6) Closure progress and plan for other Tower-A issues:

      - Solithane – Jeff said that the PO is in the buyer’s hand.  This is off-the-shelf and should be at SLAC in 1 week.  It is not hazmat.

      - Nusil silicones: the PO is completed and they are calling the vendor.  We are okay already for Tower A.  This is off-the-shelf, so Jeff expects it in a week or so. 

      - ball nests and fasteners: magnets and clips are getting epoxied in.  Fasteners are on order for next week.

      - studs and other I/F hardware: 410 studs are being dycronited (Friday); 50 hex nuts are also being dycronited; all cones and hex bushings have been fabricated; Arthur plans to pull 10 parts of each and do tests on running torque.  Then he will go to Nyloc with the new parts and will stay with them until we have a procedure that works every time.  Every stud will be run in on both ends at SLAC, using flight square and hex nuts.

      - custom socket wrench for studs: Peter is working it and will test it with Arthur

      - cone extraction tools: Peter is working it (a couple will be here next week; they are presently in the SLAC shop)

      - cable holding plate: all 18 are at SLAC; SLAC will do fit checks with the inner shipping containers and ship together the whole assembly along with the outer shipping container, in time for each tower.

      - new clips for cable holding plate: they will be in next week; 50 of each are on order

      - ESD caps: 1 week lead time, Jeff will order them off the shelf with a credit card

      - corner brackets: will be completed tomorrow but then need a complete inspection

      - clips with rounded corners and Kapton tape: same time frame; Jeff needs to order tape (check with I&T)

      - outer shipping container; will arrive at SLAC next week; then SLAC will qual test it the next week; Steve said it could be at INFN in 2 to 3 weeks, but to make it work we need manpower put on this.

      - inner shipping containers; 1 set due Oct 16th; the remainder to follow by mid November

      - base plates, including grounding screw: these are in hand, and we will add grounding screws at SLAC; 1st is due October 18th

      - cable bending tools, normal and top: Arthur said that this is in progress; the redesign is mostly finished on the old one.  The new one will be ready today or tomorrow.  They will be in the SLAC shop next week.

      - long screwdriver bit; Jeff procured assorted extensions; Robert will carry them to Pisa.  Done.

      - heat straps and tie-downs; all have been made; a set of 3 is being CMM’ed

      - cable strain relief: Jeff will call Fred Gross about what sort of foam can be used.  Done.

7) Questions from Steve:

      1.  Is it possible to use the Digi-torque for the installation of items 37 & 38 on LAT-DS-00722 rather than purchasing additional new torque screwdrivers?  Is a special torque driver needed for top corner brackets?  Answer: Pisa can use what they have.

      2.  Discussion of problem with removing the bottom row of screws Item 27 on LAT-DS-00722. Does torque value need to change? Does bit fit correctly? Are there enough bits to replace worn bits frequently enough?  The custom screw lead time is 3 to 4 weeks.  INFN will investigate these issues and report next week.  Done.

8) G&A production readiness: drawing and procedure document mods?  Martin said that drawings will be done probably Monday.  Sandro will review results on the 3 test articles tomorrow.

9) Other business:

- Closure of any remaining Tower-0 NCRs and a Tower-A assembly PRR: we need to schedule a review date around the first of November.

- Return of the EM tower: it is in the shipping container; Monday morning it will ship to SLAC

- Prepreg in Rome: it is still in the freezer in customs; SLAC is getting charged for storage; Ben wants to see the certified test values; Arthur will check with COI on this.  Ben is also interested in the earlier shipments.  Arthur will check when test results will be available.

- Sandro sent info on the CC bars.  More material would be needed Feb or March.  Alcomp is 3 to 4 months lead time.

- ESD sheets are needed in Pisa for wrapping towers.

- Foam or pads: is there any closed-cell foam that doesn’t particulate; Jeff will send information.

- Desiccant bags are needed.